the sum of three numbers equal 91. the first is five more than the second, and the third is two times the first. what are the three numbers
Answer by mananth(16946)

let the second number be x
first number = x+5
third number = 2(x+5)
...
x+x+5 +2(x+5)=91
x+x+5+2x+10=91
4x+15=91
-15
4x=76
/4
x=19
...
the numbers are 19, 24, 48
...
m.ananth@hotmail.ca